# Artificial Generational Intelligence: Cultural Accumulation in Reinforcement Learning

### v1

@@ 2.1 Partially-Observable Stochastic Games
A Partially-Observable Stochastic Game (POSG) (Kuhn, 1953) is defined by the tuple $M=\langle I,S,A,T,R,O,\gamma\rangle$. $I=\{1,\dots,N\}$ defines the set of agents, $S$ defines the set of states, and $A$ defines the direct product of action spaces for each agent, i.e., $A=A^{1}\times\dots\times A^{N}$ where $A^{i}$ is the action space for agent $i\in I$. $T$ defines the stochastic transition function, $T:S\times A\times S\to[0,1]$, $R$ defines the reward function, $R:S\times A\to\mathcal{R}$, and $O$ defines the observation function, $O:S\to[0,1]$. $\gamma\in(0,1]$ refers to the discount factor of returns. We define $\tilde{S}$ as the initial state distribution.
@@ 2.4 Generational Training
Generational training refers to training populations of agents in sequence, thus chaining together multiple learning processes (Stooke et al., 2022). We define a generation $g_{m}=\{g^{1}_{m},\dots,g^{N}_{m}\}$ as a set of $N$ agents. A new generation is generated by a process that conditions on the old generation $g_{m+1}=f(\cdot|g_{m})$. By modelling cultural accumulation over both in-context RL and RL training, $f$ could capture generational accumulation in both knowledge and skills.
@@ 3.1 Environments
Figure 2: Left: A visualization of the Goal Sequence Environment. Right: Routes travelled get shorter across generations in the TSP environment.
@@ 4 Cultural Accumulation in RL
We model cultural accumulation as taking place within POSGs. Each member of a given generation $g_{m+1}$ learns in parallel on an independent instance of the environment, which also contains static members of the previous generation, $g_{m}$. Therefore, from the perspective of a given agent $i$ in $g_{m+1}$ this is a POMDP, since the prior, static generation can be considered part of the environment. Note that in this work we focus on generational improvement and have thus introduced a separation between the development phase of an agent, where they are learning from and improving upon prior generations, and the transmission phase, where they are simply acting and being observed by the next generation. Under this formulation, we propose and investigate two distinct mechanisms for cultural accumulation in RL agents based on our in-context and in-weights dichotomy introduced above.
@@ 4.1.1 Training
To facilitate (1) during training, but not at test time, we assume access to an oracle $-i$ with fixed parameters $\hat{\theta}^{-i}$. This oracle is able to condition on the full state $s$, which we refer to as privileged information. Agent $i$ can observe the behaviours of $-i$ as they jointly act in an environment. To achieve (2), we include this representation of $-i$ in $o^{i}$ with probability $p_{\text{obs}}$ at each timestep and linearly anneal $p_{\text{obs}}$ across the $K$ trials from $1\to 0$. This ensures that the agent learns to act independently by the final trial. Perez et al. (2023) show that increasing opportunities to learn socially leads to less diversity in homogeneous populations, which could be seen as limiting independent discovery in our context, further motivating this constraint. For (3), we add some random noise to the oracle policy according to a tunable parameter $\epsilon$ [^3] For the oracle to capture sub-optimal, but goal-directed behaviour, we inject correlated noise by corrupting the privileged information in $o^{-i}$.. This encourages the agent to learn that the behaviour of other agents $-i$ in the environment may be sub-optimal, in which case $i$ should engage in selective social learning (Poulin-Dubois and Brosseau-Liard, 2016). This approach could be seen as imposing an information rate limit (Prystawski et al., 2023) between the oracle and learning agent. In Appendix A, we present this process altogether as Algorithm 2.
@@ 4.1.2 Evaluation
Having trained agents via Algorithm 2, we evaluate their in-context accumulation. During this evaluation phase, $-i$ becomes the best of the previous generation of agents $g_{m}$ [^4] This is an external selection mechanism. We investigate the emergence of selective social learning in Appendix D., replacing the oracle as we do not assume access to privileged information at test time. Agent $-i$, in transmission phase, has its internal state $\phi^{-i}$ reset between trials. A member of the new generation, in development phase, continuously updates its internal state $\phi^{i}$ across trials. The same meta-parameters $\theta^{i}$ from training are used across generations and only the internal state $\phi^{i}$ is used to update each generation. Generations are made up of $N$ agents that interact with parallel environments.
For each new generation, we store the internal state of each agent on the first step of its final trial and call it $\phi^{\prime}_{i}$. After the final trial, we then keep the previously stored internal state of the best performing agent, $\bar{\phi}=\phi^{\prime}_{i^{*}}$. We can replace the old generation, now denoted $-i$, by resetting their internal states to this new internal state, $\phi^{-i}=\bar{\phi}$ at the beginning of each trial. Thus, the next generation $g_{m+1}$ can in essence observe $g_{m}$ repeating its final trial. In practice, we mask $g_{m+1}$ from the observations $o^{-i}$ of $g_{m}$, so that $o^{-i}$ reflect the fact that agents expect not to see other agents in their final trial, as guaranteed by training attribute (2) above. Since the observations $o^{i}$ of $g_{m+1}$ can include information about the learned behaviours of $g_{m}$, we implicitly have $g_{m+1}=f(\cdot|g_{m})$, establishing a purely in-context equivalent to generational training (Section 2.4). This process is presented in Algorithm 1.
@@
3:  $s_{0}\sim\tilde{S}$ // We use a fixed initial state
4:  $\bar{\phi}=\hat{\phi}$
5:  for each generation $g\in[1,G]$ do
6:     for each population member $i\in[1,N]$ do
7:        $\phi^{i}=\hat{\phi}$
8:        $s\leftarrow s_{0}$
@@
11:              $\phi^{\prime}_{n}=\phi^{i}$ // Store agent state at beginning of last trial
@@
14:           $\phi^{-i}\leftarrow\bar{\phi}$
@@
16:              Rollout agents as in training
17:           end while
18:        end for
19:        ϕ¯¯italic-$\bar{\phi}\leftarrow$ $\phi^{\prime}_{i^{*}}$ where $i^{*}$ is the best performing agent
20:     end for
21:  end for
@@ 4.2 In-Weights Accumulation
Finally, we present an algorithm for cultural accumulation to take place over the course of training successive generations of policies (i.e., in-weights) rather than through solely updating the internal state (i.e., in-context). This is akin to generational training (Section 2.4), but without modifying the policy-gradient loss, by including the previous generation within environments on which the new generation is training. This alternative model of cultural accumulation assumes an agent’s lifetime to be an entire training run, rather than a single episode. Here, we anneal the probability of observing the previous generation $p_{\text{obs}}$ on a given environment timestep linearly over training. The corresponding approach is detailed in Algorithm 3, presented in Appendix B. In-weights RL (i.e., RL training) can be interpreted as updating procedural memory (Johnson, 2003) via skill acquisition. We thus interpret in-weights accumulation as gradual, skills-based accumulation. The results that correspond to this setting are in Section 5.2.
@@ 5 Results
For experiments in Goal Sequence, we use a Simplified Structured State Space Model (S5) (Smith et al., 2022) modified for RL (Lu et al., 2024) to encode memory, building off of the PureJaxRL codebase (Lu et al., 2022a). This model architecture runs asymptotically faster than Transformers in sequence length and outperforms RNNs in memory tasks (Morad et al., 2024). We instead simply use a GRU for Memory Sequence and TSP experiments. For Goal Sequence experiments, observations are processed by a CNN, whilst for Memory Sequence and TSP, they are processed by feed forward layers. Architectural details are provided in Appendix F and algorithmic hyperparameters in Appendix G. For all in-context experiments, a one-hot encoding of the reward received on each timestep and a one-hot encoding of the current trial number are included each agent’s observation. We provide further details on the configuration of each environment in Appendix E. All results are averaged across 10 seeds and the shaded regions of plots show standard error.
@@ 5.1 In-Context Results
As single-lifetime baselines, we use RL2 trained on: (a) episodes of equivalent length to one generation and (b) episodes of equivalent length to all cumulative generations. All in-context evaluations are performed on held-out environment instances.
Figure 3: Left: In-context accumulation on Memory Sequence. Right: Results when trained with different oracle accuracies.
@@
Figure 4: Left: In-context accumulation on Goal Sequence. Right: In-context accumulation on TSP.
@@ 5.2 In-Weights Results
TSP: We show results for in-weights accumulation in TSP in Figure 7, within Appendix C.
@@ 6 Related Work
Learning Few-Shot Learning
@@ 7 Conclusion
Limitations and Future Work: An interesting future direction would be to use learned auto-curricula (Dennis et al., 2020; Jiang et al., 2021) for deciding when agents should learn socially or independently, instead of linearly annealing the observation probability. Another extension would be to investigate cumulative culture in settings with different incentive structures, such as heterogeneous or cooperative rewards (Leibo et al., 2021; Rutherford et al., 2023). Future work could also explore learning to influence the learning process of other agents (Lu et al., 2022b, 2023) for cultural transmission. Finally, we note that whilst understanding cumulative culture has benefits for both machine learning and modelling human society, the consequences of self-improving systems should warrant consideration.
Appendix A Meta-RL Training for In-Context Accumulation
Red represents our changes to standard RL2-style training.
Algorithm 2 Training Loop for In-Context Accumulation
1:  ϕ^:=assign^italic-$\hat{\phi}:=$ init. agent state
2:  $\theta^{i}:=$ agent parameters
3:  $\hat{\theta}^{-i}:=$ oracle parameters
4:  ϵ:=assignitalic-$\epsilon:=$ oracle noise factor
5:  for train step $t\in[0,T-1]$ do
6:     $s\sim\hat{S}$
7:     $\phi\leftarrow\hat{\phi}$
8:     $B:=$ trajectory buffer
9:     for each trial $k\in[0,K-1]$ do
10:        $p_{\text{obs}}=1-k/(K-1)$
11:        while trial not done do
12:           $o\sim O(s)$
13:           if $p\sim P(p_{\text{obs}})$ then
14:              Set oracle visible in $o^{i}$
15:           end if
16:           $a^{-i}\sim\pi_{\hat{\theta}^{-i},\epsilon}(o^{-i},s)$
17:           $a^{i}\sim\pi_{\theta^{i}}(o^{i}|\phi^{i})$
18:           $\phi^{i}\leftarrow f_{\theta^{i}}(o^{i}|\phi^{i})$
19:           $s,r\sim T(s,a)$
20:           $B\leftarrow(o^{i},a^{i},r^{i})$ // Append transition to buffer
21:        end while
22:     end for
23:     $\theta^{i}\leftarrow$ update$(\theta^{i},B)$
24:  end for
Appendix B In-Weights Accumulation
@@
2:  $\bar{\theta}\sim\hat{\theta}$ // Initialise the “random” first generation parameters
3:  for each generation $g\in[1,G]$ do
4:     for each population member $n\in[1,N]$ do
@@
11:              $o\sim O(s)$
12:              if $p\sim P(p_{\text{obs}})$ then
13:                 Set oracle visible in $o^{i}$
@@
18:              $B\leftarrow(o^{i},a^{i},r^{i})$ // Append transition to buffer
@@
20:           $\theta_{n}\leftarrow\text{update}(\theta_{n},B)$
@@
22:        $\bar{\theta}\leftarrow\theta_{n^{*}}\text{where $n^{*}$ is the best performingagent}$
@@
Appendix C In-Weights Accumulation in TSP

### v2

@@ 2.1 Partially-Observable Stochastic Games
A Partially-Observable Stochastic Game (POSG) (Kuhn, 1953) is defined by the tuple $M=\langle I,S,A,T,R,O,\gamma\rangle$. $I=\{1,\dots,N\}$ defines the set of agents, $S$ defines the set of states, and $A$ defines the direct product of action spaces for each agent, i.e., $A=A^{1}\times\dots\times A^{N}$ where $A^{i}$ is the action space for agent $i\in I$. $T$ defines the stochastic transition function, $T:S\times A\times S\to[0,1]$, $R$ defines the reward function, $R:S\times A\to\mathcal{R}$, and $O$ defines the observation function, $O:\mathcal{O}\times S\to[0,1]$ where $\mathcal{O}$ is the observation space. $\gamma\in(0,1]$ refers to the discount factor of returns.
@@ 2.4 Generational Training
Generational training refers to training populations of agents in sequence, thus chaining together multiple learning processes (Stooke et al., 2022). We define a generation $g_{m}=\{g^{1}_{m},\dots,g^{N_{\text{pop}}}_{m}\}$ as a population of $N_{\text{pop}}$ agents. Each $g^{n}_{m}$ can therefore be model weights, or an internal state in the case of in-context RL. We distinguish $N_{\text{pop}}$ from the $N$ used in Section 2.1, because each member of a generation of size $N_{\text{pop}}$ may be within its own POSG of $N$ agents. We use $n$ to index a specific member of a generation. A new generation is generated by a process that conditions on the old generation $g_{m+1}=f(g_{m})$. By modelling cultural accumulation over both in-context RL and RL training, $f$ could capture generational accumulation in both knowledge and skills.
@@ 3.1 Environments
Figure 2: Left: A visualization of the Goal Sequence Environment. Right: Routes travelled get shorter across generations in the TSP environment. Visualisation implementation is based on Jumanji (Bonnet et al., 2024).
@@ 4 Cultural Accumulation in RL
We model cultural accumulation as taking place within POSGs. Each member of a given generation $g_{m+1}$ learns in parallel on an independent instance of the environment, which also contains static members of the previous generation, $g_{m}$. Therefore, from the perspective of a given agent $n$ in $g_{m+1}$ this is a POMDP, since the prior, static generation can be considered part of the environment. In this work, we focus on generational improvement and have thus introduced a separation between the development phase of an agent, where they are learning from and improving upon prior generations, and the transmission phase, where they are simply acting and being observed by the next generation. Under this formulation, we propose and investigate two distinct mechanisms for cultural accumulation in RL agents based on our in-context and in-weights dichotomy introduced in Section 3.
@@ 4.1.1 Training
Algorithm 1 Training Loop for In-Context Accumulation (changes to RL2 in red)
1:  ϕ^:=assign^italic-$\hat{\phi}:=$ init. agent state
2:  $\hat{\theta}:=$ init. agent parameters
3:  $\theta^{\texttt{O}}:=$ oracle parameters
4:  ϵ:=assignitalic-$\epsilon:=$ oracle noise factor
5:  $\theta^{n}\leftarrow\hat{\theta}$ // Initialise the agent parameters
6:  for train step $t\in[0,T-1]$ do
7:     $s\sim\hat{S}$
8:     $\phi^{n}\leftarrow\hat{\phi}$
9:     $B:=$ trajectory buffer
10:     for each trial $k\in[0,K-1]$ do
11:        $p_{\text{obs}}=1-k/(K-1)$
12:        while trial not done do
13:           $o^{n}\sim O(s)$
14:           if $\texttt{IsVisible}\sim\text{Bernoulli}(p_{\text{obs}})$ then
15:              Set oracle visible in $o^{n}$
16:           end if
17:           $a^{\texttt{O}}\sim\pi_{\theta^{\texttt{O}},\epsilon}(s)$, $a^{n}\sim\pi_{\theta^{n}}(o^{n}|\phi^{n})$
18:           $\phi^{n}\leftarrow f_{\theta^{n}}(o^{n}|\phi^{n})$
19:           $s,r\sim T(s,a)$
20:           $B\leftarrow(o^{n},a^{n},r^{n})$ // Append transition to buffer
21:        end while
22:     end for
23:     $\theta^{n}\leftarrow$ update$(\theta^{n},B)$
24:  end for
@@
To facilitate (1) during training, but not at test time, we assume access to an oracle O with fixed parameters $\hat{\theta}^{\texttt{O}}$. O is able to condition on the full state $s$, which we refer to as privileged information. Agent $n$ can observe the behaviours of O as they jointly act in an environment. To achieve (2), we include this representation of O in $o^{n}$ with probability $p_{\text{obs}}$ at each timestep and linearly anneal $p_{\text{obs}}$ across the $K$ trials from $1\to 0$. This ensures that the agent learns to act independently by the final trial. Perez et al. (2023) show that increasing opportunities to learn socially leads to less diversity in homogeneous populations, which could be seen as limiting independent discovery in our context, further motivating this constraint. For (3), we add some random noise to the oracle policy according to a tunable parameter $\epsilon$ [^3] For the oracle to capture sub-optimal, but goal-directed behaviour, we inject correlated noise by corrupting the privileged information in $o^{\texttt{O}}$.. This encourages the agent to learn that the behaviour of other agents O in the environment may be sub-optimal, in which case agent $n$ should engage in selective social learning (Poulin-Dubois and Brosseau-Liard, 2016). This approach could be seen as imposing an information rate limit (Prystawski et al., 2023) between the oracle and learning agent. We present this process altogether as Algorithm 1, with changes to standard RL2 in red.
@@ 4.1.2 Evaluation
3:  $s_{0}\sim\hat{S}$ // Sample initial state
4:  $\tilde{\phi}^{n^{*}}_{0}\leftarrow\hat{\phi}$ // Define an initial reset state
5:  for each generation $m\in[1,M]$ do
6:     for each population member $n\in[1,N_{\text{pop}}]$ do
7:        $\phi^{n}_{m}\leftarrow\hat{\phi}$
8:        $s^{n}\leftarrow s_{0}$
@@
11:              $\tilde{\phi}^{n}_{m}\leftarrow\phi^{n}_{m}$ // Store agent state at beginning of last trial as reset state
@@
14:           $\phi^{n}_{m-1}\leftarrow\tilde{\phi}^{n^{*}}_{m-1}$ // Set previous generation state to best reset state from that generation
@@
16:              $o^{n}_{m-1},o^{n}_{m}\sim O(s)$
17:              if $\texttt{IsVisible}\sim\text{Bernoulli}(p_{\text{obs}})$ then
18:                 Set previous generation visible in $o^{n}_{m}$
19:              end if
20:              $a^{n}_{m-1}\sim\pi_{\theta}(o^{n}_{m-1}|\phi^{n}_{m-1}),\,a^{n}_{m}\sim\pi_{\theta}(o^{n}_{m}|\phi^{n}_{m})$
21:              $\phi^{n}_{m-1}\leftarrow f_{\theta}(o^{n}_{m-1}|\phi^{n}_{m-1}),\,\phi^{n}_{m}\leftarrow f_{\theta}(o^{n}_{m}|\phi^{n}_{m})$
22:              $s^{n},r^{n}\sim T(s^{n},a^{n})$
23:           end while
24:        end for
25:        $n^{*}=\text{argmax}_{n\in[1,N_{\text{pop}}]}\sum_{t}\mathbb{I}[k=K-1]r^{n}_{m}$ // Select best population member based on final trial performance
26:     end for
27:  end for
Having trained agents via Algorithm 1, we evaluate their in-context accumulation. During this evaluation phase, O is replaced by the best [^4] This is an external selection mechanism. We investigate the emergence of selective social learning in Appendix C. of the previous generation $n^{*}$, as we do not assume access to privileged information at test time. This previous best agent, now in transmission phase, has its internal state reset between trials. A member of the new generation, in development phase, continuously updates its internal state across trials. The same meta-parameters $\theta$ from training are used across generations and only the internal states $\phi$ are used to update each generation. Generations are made up of $N$ agents that interact with $N$ independent environments.
For each generation $g_{m}$, we store the internal state of each agent on the first step of its final trial and call this the “reset state” $\tilde{\phi}^{n}_{m}$. After the final trial, we then keep the previously stored reset state of the best performing agent $\tilde{\phi}^{n^{*}}_{m}$ in terms of rewards received in its last trial. When introducing the next generation $g_{m+1}$, generation $g_{m}$ transitions into transmission phase and we represent the adapted behaviour of the best performing agent by resetting the internal states $\phi^{1:N}_{m}\leftarrow\tilde{\phi}^{n^{*}}_{m}$ at the beginning of each trial. Thus, the next generation $g_{m+1}$ can in essence observe the best of $g_{m}$ repeating its final trial. In practice, we mask $g_{m+1}$ from the observations of $g_{m}$ in order to reflect the fact that agents do not expect to see other agents in their final trial, as guaranteed by training attribute (2) above. Since the observations of $g_{m+1}$ can include information about the learned behaviours of $g_{m}$, we implicitly have $g_{m+1}=f(g_{m})$, establishing a purely in-context equivalent to generational training (Section 2.4). This process is presented in Algorithm 2.
@@ 4.2 In-Weights Accumulation
Finally, we present an algorithm for cultural accumulation to take place over the course of training successive generations of policies (i.e., in-weights) rather than through solely updating the internal state (i.e., in-context). This is akin to generational training (Section 2.4), but without modifying the policy-gradient loss, by simply including the previous generation within environments on which the new generation is training, so that their actions can be observed and learned from. This alternative model of cultural accumulation assumes an agent’s lifetime to be an entire training run, rather than a single episode. Here, we anneal the probability of observing the previous generation $p_{\text{obs}}$ on a given environment timestep linearly over training. The corresponding approach is detailed in Algorithm 3, presented in Appendix A. In-weights RL (i.e., RL training) can be interpreted as updating procedural memory (Johnson, 2003) via skill acquisition. We thus interpret in-weights accumulation as gradual, skills-based accumulation. Note that there is no separation between "training" and "evaluation" in this case, because the in-weights algorithm views cultural accumulation as part of the training. This is in contrast to in-context accumulation, which is an evaluation-time algorithm executed by appropriately trained meta-policies. The results that correspond to this setting are in Section 5.2.
@@ 5 Results
Figure 3: Left: In-context accumulation during evaluation on Memory Sequence. Right: Evaluation results following training with different oracle accuracies.
For each of our experiments, we use a Simplified Structured State Space Model (S5) (Smith et al., 2022) modified for RL (Lu et al., 2024) to encode memory, building on the PureJaxRL codebase (Lu et al., 2022a). This model architecture runs asymptotically faster than Transformers in sequence length and outperforms RNNs in memory tasks (Morad et al., 2024). PPO (Schulman et al., 2017) is used as the RL training algorithm. For Goal Sequence experiments, observations are processed by a CNN, whilst for Memory Sequence and TSP, they are processed by feed forward layers. Architectural details are provided in Appendix E and algorithmic hyperparameters in Appendix F. For all in-context experiments, a one-hot encoding of the reward received on each timestep and a one-hot encoding of the current trial number are included each agent’s observation. We provide further details on the configuration of each environment in Appendix D. All results are averaged across 10 seeds and the shaded regions of plots show standard error.
Figure 4: Left: In-context accumulation during evaluation on Goal Sequence. Right: In-context accumulation during evaluation on TSP.
@@ 5.1 In-Context Results
As single-lifetime baselines, we use RL2 trained on: (a) episodes of equivalent length to one generation and (b) episodes of equivalent length to all cumulative generations. All in-context evaluations are performed on held-out environment instances for the same number of trials.
@@ 5.2 In-Weights Results
TSP: We show results for in-weights accumulation in TSP in Figure 7, within Appendix B.
@@ 6 Related Work
Learning Few-Shot Imitation
@@ 7 Conclusion
Limitations and Future Work: An interesting future direction would be to use learned auto-curricula (Dennis et al., 2020; Jiang et al., 2021) for deciding when agents should learn socially or independently, instead of linearly annealing the observation probability. Another extension would be to investigate cumulative culture in settings with different incentive structures, such as heterogeneous or cooperative rewards (Rutherford et al., 2023). Future work could also study ways we can evolve the process of cultural accumulation itself (Lu et al., 2023a) or learn to influence the learning process of other agents (Lu et al., 2022b, 2023b) for cultural transmission. Finally, we note that whilst understanding cumulative culture has benefits for both machine learning and modelling human society, the consequences of self-improving systems should warrant consideration.
Acknowledgments
Jonathan Cook is supported by the ESPRC Centre for Doctoral Training in Autonomous Intelligence Machines and Systems EP/S024050/1. Jakob Foerster is partially funded by the UKI grant EP/Y028481/1 (originally selected for funding by the ERC). Jakob Foerster is also supported by the JPMC Research Award and the Amazon Research Award.
Appendix A In-Weights Accumulation
@@
2:  $\tilde{\theta}^{n^{*}}_{0}\leftarrow\hat{\theta}$ // Initialise the “random” first generation parameters
3:  for each generation $m\in[1,G]$ do
4:     for each population member $n\in[1,N_{\text{pop}}]$ do
@@
11:              $o^{n}_{m-1},o^{n}_{m}\sim O(s)$
12:              if $p\sim\text{Bernoulli}(p_{\text{obs}})$ then
13:                 Set previous generation visible in $o^{n}_{m}$
@@
18:              $B\leftarrow(o^{n}_{m},a^{n}_{m},r^{n}_{m})$ // Append transition to buffer
@@
20:           $\theta^{n}_{m}\leftarrow\text{update}(\theta^{n}_{m},B)$
@@
22:        $n^{*}=\text{argmax}_{n\in[1,N_{\text{pop}}]}\sum_{t=T-1}r^{n}_{m}$ // Select best population member based on final episode performance
@@
Appendix B In-Weights Accumulation in TSP